The inventory turnover for an industry is 6 (every two months) but Slow Corp. turns over its inventory 4 times a year (every three months). If annual sales are $1,000,000 and the interest cost to carry inventory is 12 percent, what is the potential savings in interest expense if the firm achieves the industry for the turnover of its inventory?
Answer #1

Industry turnover ratio = Sales / Avg inventory

6 = 1000000 / Avg inventory

Avg inventory = 1000000 / 6 = 166666.67

Slow corp avg inventory = 1000000 / 4 = 250000

To achieve industry standard, avg inventory has to reduce by = 250000 - 166666.67 = 83333.33

saving in interest expense = 83333.33 * 0.12 = 10000 (Answer)